# Optimal Computation in Anonymous Dynamic Networks

### v8

@@ 4.2 Stabilizing Algorithm for Leaderless Networks
Our local algorithm is as follows. Each agent $p$ runs Listing 1 on its own vista $\mathcal{V}$, obtaining a pair $(t,S)$. If $t=-1$ or $S$ is not a homogeneous system of $k-1$ independent linear equations in $k$ variables, then $p$ outputs $\{(\lambda(p),1)\}$, where $\lambda(p)$ is its own input. Otherwise, since the rank of the coefficient matrix of $S$ is $k-1$, the general solution to $S$ has exactly one free parameter, due to the Rouché–Capelli theorem. Therefore, by Gaussian elimination, it is possible to express every variable $x_{i}$ as a rational multiple of $x_{1}$, i.e., $x_{i}=\alpha_{i}x_{1}$ for some $\alpha_{i}\in\mathbb{Q}^{+}$ (recall that the coefficients of $S$ are integers). Let $L_{t}=\{w_{1},w_{2},\dots,w_{k}\}$ and $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$. For every node $v_{i}\in L_{0}$, define $\beta_{i}\in\mathbb{Q}^{+}$ as $\beta_{i}=\sum_{w_{j}\in L_{t}\text{ descendant of }v_{i}}\alpha_{j}$, and let $\beta=\sum_{i}\beta_{i}$. Then, $p$ outputs
@@
The correctness and stabilization time of the above algorithm directly follow from Lemma 4.2. Specifically, if not all agents have the same input (hence $n\geq c\geq 2$), then at any round $t^{\prime}\geq\tau(2n-2)\geq\tau(2n-3)+1\geq\tau(2n-c-1)+1$, Condition 2 of Lemma 4.2 is met, and the system $S$ is satisfied by the anonymities of the nodes in $L_{t}$. Thus, $a(v_{i})=\alpha_{i}a(v_{1})$ for all $v_{i}\in L_{0}$, and therefore $\beta_{i}/\beta=a(v_{i})/n$. We conclude that, for any input assignment $\lambda$, the algorithm stabilizes on the correct output $\frac{1}{n}\cdot\mu_{\lambda}$ within $\tau(2n-2)$ rounds.
In the special case where all agents have the same input, then the default output $\{(\lambda(p),1)\}$ is correct since round 00. Also, whenever $S$ is a homogeneous system of $k-1$ independent linear equations, the algorithm still returns the correct output $\{(\lambda(p),1)\}$, because all nodes in level $L_{0}$ have the same label $\lambda(p)$. Hence, in this particular case the output is always correct. ∎
@@ 4.3 Terminating Algorithm for Leaderless Networks
For every $\tau\geq 1$ and $N\geq 1$, there is an algorithm that computes the Input Frequency function $F_{IF}$ on all networks in $\bigcup_{n\leq N}\mathcal{N}^{\tau}_{n,0}$ and terminates in at most $\tau(n+N-2)+1$ rounds.
@@
Given $\tau$ and $N$, our terminating local algorithm is as follows. Run Listing 1 on the agent’s vista $\mathcal{V}$, obtaining a pair $(t,S)$, and then do the same computations as in the algorithm for Theorem 4.4. If $t\geq 0$ and the current round $t^{\prime}$ satisfies $t^{\prime}\geq t+\tau(N-1)+1$, then the output is correct, and the agent terminates.
The correctness of this algorithm is a direct consequence of Lemma 4.2. Indeed, the algorithm only terminates when $t\geq 0$ and $t^{\prime}\geq t+\tau(N-1)+1\geq t+\tau(n-1)+1$, and hence it gives the correct output because Condition 1 of Lemma 4.2 is met. As for the running time, assume that $t^{\prime}=\tau(n+N-2)+1\geq\tau(2n-c-1)+1$. Since Condition 2 of Lemma 4.2 is met, we have $0\leq t\leq\tau(n-c)\leq\tau(n-1)$. Thus, $t^{\prime}=\tau(n+N-2)+1\geq t+\tau(N-1)+1$, and the algorithm terminates at round $t^{\prime}$. ∎
@@
For every $d\geq 1$, there is an algorithm that computes the Input Frequency function $F_{IF}$ on all networks in $\bigcup_{n\geq 1}\mathcal{N}^{[d]}_{n,0}$ and terminates in at most $\tau(n-1)+d+1\leq dn+1$ rounds, where $\tau$ is the dynamic disconnectivity of the network.
@@
We can repeat the proof of Theorem 4.5 verbatim, replacing the termination condition with $t^{\prime}\geq t+d+1$ and invoking Corollary 4.3 in lieu of Lemma 4.2. ∎
Note that knowledge of an upper bound $D\geq d$ is actually enough for the algorithm in Corollary 4.6 to work. Indeed, with the termination condition $t^{\prime}\geq t+D+1$, the same algorithm computes $F_{IF}$ on all networks in $\bigcup_{d\leq D}\bigcup_{n\geq 1}\mathcal{N}^{[d]}_{n,0}$ and terminates in at most $\tau(n-1)+D+1\leq d(n-1)+D+1$ rounds.
@@ 4.4 Stabilizing Algorithm for Networks with Leaders
The correctness follows from the fact that, as shown in Theorem 4.4, at any round $t^{\prime}\geq\tau(2n-c-1)+1$ we have $\beta_{i}/\beta=a(v_{i})/n$ for all $1\leq i\leq k^{\prime}$. Adding up these equations for all $i\in\{j_{1},j_{2},\dots,j_{l}\}$, we obtain $\beta^{\prime}/\beta=\ell/n$, and therefore $n=\ell\beta/\beta^{\prime}$. We conclude that
@@
Thus, within $\tau(2n-c-1)+1$ rounds, the algorithm stably outputs the anonymities of all nodes in $L_{0}$. As observed in Section 2, this is equivalent to computing the Input Multiset function $F_{IM}$.
If there are both leader and non-leader agents in the system, there are $c\geq 2$ distinct inputs, and therefore the stabilization time is at most $\tau(2n-2)\geq\tau(2n-c-1)+1$ rounds, as desired. In the special case where all agents are leaders, we have $\ell=n$, and the algorithm returns the correct output at every round since round 00. Thus, the stabilization time is $\tau(2n-2)$ rounds in any case. ∎
@@ 5 Terminating Algorithm for Networks with Leaders
Secondly, networks with more than one leader significantly add to the difficulty of the problem. For instance, once a terminating algorithm for networks with a unique leader has been designed, one may be tempted to simply adapt it to the multi-leader case by setting the anonymity of the leader node in the history tree to $\ell>1$ instead of $1$. However, this approach would be inconsistent. Indeed, while the history tree of a network with a unique leader contains a single leader branch, all of whose nodes have anonymity $1$, this may not be the case for networks with multiple leaders. In such a network, as soon as some leaders get disambiguated, the leader node branches into several children nodes whose anonymities are unknown (we only know that their sum is $\ell$). Moreover, some leader nodes may be missing from the vista of other leaders. Thus, in the case of multi-leader networks, we must deal with the fact that even the vista of a leader may have levels where the sum of the anonymities of any subset of nodes is unknown.
@@
In Section 5.2, we approach the general multi-leader problem by repeatedly guessing the anonymity of a leader node in the history tree and invoking the subroutine of Section 5.3 to confirm our guesses. This procedure introduces a further bottleneck that causes our final algorithm to have a running time of $O(\tau\ell^{2}n)$ rounds.
@@ 5.1 Naive Termination Strategies Are Incorrect
Figure 4: An example of a dynamic network where the naive technique of Section 4.4 fails to provide a correct termination condition. The white nodes in the history tree are not in the history of the leader at the last round; the red edges not in the vista are not drawn. Same-colored agents have equal inputs. For the first four rounds, from the leader’s perspective, this network is indistinguishable from the complete bipartite graph $K_{1,4}$. Thus, throughout this time, the leader is unaware of the agents labeled $p_{3}$, and therefore cannot compute the total number of agents.
@@
Unfortunately, this strategy fails on the network in Figure 5. Here the leader’s vista at levels $L_{0}$ and $L_{1}$ causes the algorithm to count only $n^{\prime}=5$ agents (i.e., one leader, two agents represented by the purple node, and one agent represented by the yellow node). Afterwards, the leader has to wait until round $k-5$ for the appearance of the node that was missing from $L_{1}$. Since $k$ is arbitrary, this type of strategy is bound to fail no matter what the waiting time is.
@@ 5.2 Main Algorithm
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 7 Assign $t^{\ast}:=-1$
8 For $x:=\ell$ downto $1$
9 Assign $(n^{\prime},t):=\text{ApproxCount}(\mathcal{V},s,x,\ell)$ see Listing 3 in Section 5.3
@@ 17 Assign $s:=t^{\ast}+1$
[…]
@@
[…]
Observe that, since there can be at most $\ell-1$ branching leader nodes in $\mathcal{V}$, at least one approximation $n^{\prime}>0$ of $n$ is computed within the $\ell$th iteration of the loop at Lines 6–17. This occurs within $t^{\ast}\leq\ell\Delta=t^{\prime}-n$ levels. Because all nodes in these levels must appear in $\mathcal{V}$, the condition $a(\vartheta)=a(\vartheta^{\prime})$ is satisfied in all intervals $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, …, $\mathcal{L}_{\ell-b+1}$. Due to Lemma 5.1 (ii), we conclude that all such intervals must yield the correct approximation $n^{\prime}=n$. So, every time Line 15 is executed, we have $n^{\ast}=n^{\prime}$, and the algorithm does not return "Unknown". Finally, when Line 19 is reached, we have $t^{\ast}\leq t^{\prime}-n=t^{\prime}-n^{\ast}$, and therefore "Unknown" is not returned. Thus, an output other than "Unknown" is returned within the desired number of rounds. ∎

### v9

@@ 4.2 Stabilizing Algorithm for Leaderless Networks
Our local algorithm is as follows. Each agent $p$ runs Listing 1 on its own vista $\mathcal{V}$, obtaining a pair $(t,S)$. If t=1t=-1 or $S$ is not a homogeneous system of k1k-1 independent linear equations in $k$ variables, then $p$ outputs $\{(\lambda(p),1)\}$, where $\lambda(p)$ is its own input. Otherwise, since the rank of the coefficient matrix of $S$ is k1k-1, the general solution to $S$ has exactly one free parameter, due to the Rouché–Capelli theorem. Therefore, by Gaussian elimination, it is possible to express every variable $x_{i}$ as a rational multiple of $x_{1}$, i.e., $x_{i}=\alpha_{i}x_{1}$ for some $\alpha_{i}\in\mathbb{Q}^{+}$ (recall that the coefficients of $S$ are integers). Let $L_{t}=\{w_{1},w_{2},\dots,w_{k}\}$ and $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$. For every node $v_{i}\in L_{0}$, define $\beta_{i}\in\mathbb{Q}^{+}$ as $\beta_{i}=\sum_{w_{j}\in L_{t}\text{ descendant of }v_{i}}\alpha_{j}$, and let $\beta=\sum_{i}\beta_{i}$. Then, $p$ outputs
@@
The correctness and stabilization time of the above algorithm directly follow from Lemma 4.2. Specifically, if not all agents have the same input (hence $n\geq c\geq 2$), then, at any round $t^{\prime}$ such that $t^{\prime}\geq\tau(2n-2)\geq\tau(2n-c)$, Condition 2 of Lemma 4.2 is met, and the system $S$ is satisfied by the anonymities of the nodes in $L_{t}$. Thus, $a(v_{i})=\alpha_{i}a(v_{1})$ for all $v_{i}\in L_{0}$, and therefore $\beta_{i}/\beta=a(v_{i})/n$. We conclude that, for any input assignment $\lambda$, the algorithm stabilizes on the correct output $\frac{1}{n}\cdot\mu_{\lambda}$ within $\tau(2n-2)$ rounds.
In the special case where all agents have the same input, the default output $\{(\lambda(p),1)\}$ is correct since round 0. Also, whenever $S$ is a homogeneous system of k1k-1 independent linear equations, the algorithm still returns the correct output $\{(\lambda(p),1)\}$, because all nodes in level $L_{0}$ have the same label $\lambda(p)$. Hence, in this particular case the output is always correct. ∎
@@ 4.3 Terminating Algorithm for Leaderless Networks
For every $\tau\geq 1$ and $N\geq 1$, there is an algorithm that computes the Input Frequency function $F_{IF}$ on all networks in $\bigcup_{n\leq N}\mathcal{N}^{\tau}_{n,0}$ and terminates in at most $\tau(n+N-2)$ rounds.
@@
Given $\tau$ and $N$, our terminating local algorithm is as follows. Run Listing 1 on the agent’s vista $\mathcal{V}$, obtaining a pair $(t,S)$, and then do the same computations as in the algorithm for Theorem 4.4. If $t\geq 0$ and the current round $t^{\prime}$ satisfies $t^{\prime}\geq t+\tau(N-1)+1$, then the output is certifiably correct, and the agent terminates. Moreover, if $t^{\prime}=\tau(N-1)$ and the level $L_{0}$ of $\mathcal{V}$ contains a single node, then the agent outputs $\{(\lambda(p),1)\}$ and terminates, where $\lambda(p)$ is the agent’s own input.
Due to Corollary 3.4, the latter condition is verified only if all agents have the same input (because level $L_{0}$ of $\mathcal{V}$ must be complete at round $\tau(N-1)\geq\tau(n-1)$). In this case, it is safe to conclude that the correct output is $\{(\lambda(p),1)\}$, and termination occurs within $\tau(N-1)\leq\tau(n+N-2)$ rounds, as desired (obviously, $n\geq 1$ holds).
In all other cases, the correctness of the algorithm is a direct consequence of Lemma 4.2. Indeed, if the algorithm terminates when $t\geq 0$ and $t^{\prime}\geq t+\tau(N-1)+1\geq t+\tau(n-1)+1$, it gives the correct output because Condition 1 of Lemma 4.2 is met. Thus, the algorithm cannot terminate with an incorrect output.
As for the running time, we may leverage Lemma 4.2 with $c\neq 1$, since we have already shown that for $c=1$ termination occurs within $\tau(n+N-2)$ rounds. So, assume that $c\geq 2$ and $t^{\prime}=\tau(n+N-2)\geq\tau(2n-2)\geq\tau(2n-c)$. Since Condition 2 of Lemma 4.2 is met, we have $0\leq t\leq\tau(n-c+1)-1\leq\tau(n-1)-1$. Thus, $t^{\prime}=\tau(n+N-2)\geq t+\tau(N-1)+1$, and the algorithm terminates at round $t^{\prime}$, as desired. ∎
@@
For every $d\geq 1$, there is an algorithm that computes the Input Frequency function $F_{IF}$ on all networks in $\bigcup_{n\geq 1}\mathcal{N}^{[d]}_{n,0}$ and terminates in at most $\tau(n-1)+d\leq\tau(2n-2)$ rounds, where $\tau$ is the dynamic disconnectivity of the network.
@@
We can repeat the proof of Theorem 4.5 verbatim, replacing the termination condition $t^{\prime}\geq t+\tau(N-1)+1$ with $t^{\prime}\geq t+d+1$ and the condition $t^{\prime}=\tau(N-1)$ with $t^{\prime}=d$. Then Corollary 4.3 is invoked in lieu of Lemma 4.2, recalling that $d\leq\tau(n-1)$. ∎
Note that knowledge of an upper bound $D\geq d$ is actually enough for the algorithm in Corollary 4.6 to work. Indeed, with the termination condition $t^{\prime}\geq t+D+1$, the same algorithm computes $F_{IF}$ on all networks in $\bigcup_{d\leq D}\bigcup_{n\geq 1}\mathcal{N}^{[d]}_{n,0}$ and terminates in at most $\tau(n-1)+D$ rounds.
@@ 4.4 Stabilizing Algorithm for Networks with Leaders
The correctness follows from the fact that, as shown in Theorem 4.4, at any round $t^{\prime}\geq\tau(2n-c)$ we have $\beta_{i}/\beta=a(v_{i})/n$ for all $1\leq i\leq k^{\prime}$. Adding up these equations for all $i\in\{j_{1},j_{2},\dots,j_{l}\}$, we obtain $\beta^{\prime}/\beta=\ell/n$, and therefore $n=\ell\beta/\beta^{\prime}$. We conclude that
@@
Thus, within $\tau(2n-c)$ rounds, the algorithm stably outputs the anonymities of all nodes in $L_{0}$. As observed in Section 2, this is equivalent to computing the Input Multiset function $F_{IM}$.
If there are both leader and non-leader agents in the system, there are $c\geq 2$ distinct inputs, and therefore the stabilization time is at most $\tau(2n-2)\geq\tau(2n-c)$ rounds, as desired. In the special case where all agents are leaders, we have $\ell=n$, and the algorithm returns the correct output at every round since round 0. Thus, the stabilization time is $\tau(2n-2)$ rounds in any case. ∎
@@ 5 Terminating Algorithm for Networks with Leaders
Secondly, networks with more than one leader significantly add to the difficulty of the problem. For instance, once a terminating algorithm for networks with a unique leader has been designed, one may be tempted to simply adapt it to the multi-leader case by setting the anonymity of the leader node in the history tree to $\ell>1$ instead of $1$. Unfortunately, this approach neglects one important factor. Indeed, while the history tree of a network with a unique leader contains a single leader branch, all of whose nodes have anonymity $1$, this may not be the case in a network with multiple leaders. In such a network, as soon as some leaders get disambiguated, the leader node branches into several children nodes whose anonymities are unknown (we only know that their sum is $\ell$). Moreover, some leader branches may be missing from the vistas of other leaders for several rounds. Thus, in the case of multi-leader networks, we must deal with the fact that even the vista of a leader may have levels where the sum of the anonymities of any subset of nodes is unknown.
@@
In Section 5.2, we approach the general multi-leader problem by repeatedly guessing the anonymity of a leader node in the history tree and invoking the subroutine of Section 5.3 to confirm our guesses. This procedure introduces additional overhead, bringing the running time of our final algorithm to $O(\tau\ell^{2}n)$ rounds.
@@ 5.1 Naive Termination Strategies Are Incorrect
Figure 4: An example of a dynamic network where the naive technique of Section 4.4 fails to provide a correct termination condition. The white nodes in the history tree are not in the vista of the leader at the last round; the red edges not in the vista are not drawn. Same-colored agents have equal inputs. For the first four rounds, from the leader’s perspective, this network is indistinguishable from the complete bipartite graph $K_{1,4}$. Thus, throughout this time, the leader is unaware of the agents labeled $p_{3}$, and therefore cannot compute the total number of agents.
@@
Unfortunately, this strategy fails on the network in Figure 5. Here the leader’s vista at levels $L_{0}$ and $L_{1}$ causes the algorithm to count only $n^{\prime}=5$ agents (i.e., one leader, two agents represented by the purple node, and one agent represented by the yellow node). Afterwards, the leader has to wait until round k5k-5 for the appearance of the node that was missing from $L_{1}$. Since $k$ is arbitrary, this type of strategy is bound to fail no matter what the waiting time is.
@@ 5.2 Main Algorithm
Discrepancy $\delta$. Suppose that ApproxCount is invoked with arguments $\mathcal{V}$, $s$, $x$, $\ell$, where $1\leq x\leq\ell$, and let $\vartheta$ be the first leader node in level $L_{s}$ of $\mathcal{V}$ (if $\vartheta$ does not exist, the procedure immediately returns the error message $n^{\prime}=\texttt{"MissingNodes"}$). We define the discrepancy $\delta$ as the ratio $x/a(\vartheta)$. Clearly, $\delta\leq\ell$. Note that, since $a(\vartheta)$ is not a-priori known by the agent executing ApproxCount, then neither is $\delta$.
@@
Error messages. If $L_{s}$ contains no leader nodes, the procedure returns the error message $n^{\prime}=\texttt{"MissingNodes"}$. If, before gathering enough information on $n$, the procedure encounters a descendant of $\vartheta$ with more than one child in $\mathcal{V}$, it returns the error message $n^{\prime}=\texttt{"StrandTooShort"}$. If it determines that the conditional anonymity of a node is not an integer, it returns the error message $n^{\prime}=\texttt{"WrongGuess"}$. Finally, if it determines that the sum $\ell^{\prime}$ of the conditional anonymities of the leader nodes is not $\ell$, it returns $n^{\prime}=\texttt{"MissingNodes"}$ if $\ell^{\prime}<\ell$ and $n^{\prime}=\texttt{"WrongGuess"}$ if $\ell^{\prime}>\ell$.
@@
    If $x=a(\vartheta)=a(\vartheta^{\prime})$, then $n^{\prime}\neq\texttt{"WrongGuess"}$.
@@
    If $t^{\prime}\geq s+(\ell+2)(n-1)$, then $s\leq t\leq s+(\ell+1)(n-1)$ and $n^{\prime}\neq\texttt{"MissingNodes"}$. Moreover, if $n^{\prime}=\texttt{"StrandTooShort"}$, then $L_{t}$ contains a leader node with at least two children in $\mathcal{V}$.∎
@@ 7 Assign $t^{\ast}:=-1$
8 For $x:=\ell$ downto 11
9 Assign $(n^{\prime},t):=\text{ApproxCount}(\mathcal{V},s,x,\ell)$ # see Listing 3 in Section 5.3
@@ 17 Assign $s:=t^{\ast}+1$
Let $(n^{\prime},t)$ be the pair of values returned by ApproxCount. If $n^{\prime}=\texttt{"MissingNodes"}$, this is evidence that $\mathcal{V}$ is still missing some relevant nodes, and therefore "Unknown" is immediately returned (Line 11). If $n^{\prime}=\texttt{"StrandTooShort"}$, then a descendant of $\vartheta$ with multiple children in $\mathcal{V}$ was found, say at level $L_{t}$, before an approximation of $n$ could be determined. As this is an undesirable event, the algorithm moves $\mathcal{L}_{i}$ after $L_{t}$ and tries again to estimate $n$ (Line 12). If $n^{\prime}=\texttt{"WrongGuess"}$, then $x$ may not be the correct anonymity of the leader node $\vartheta$ (see the description of ApproxCount), and therefore the algorithm calls ApproxCount again with the same starting level $L_{s}$, but now with $x=\ell-1$. If $n^{\prime}=\texttt{"WrongGuess"}$ is returned again, then $x=\ell-2$ is tried, and so on. After all possible assignments down to $x=1$ have failed, the algorithm just moves $\mathcal{L}_{i}$ forward and tries again from $x=\ell$.
@@
Correctness and running time. We will prove that, if the output of Listing 2 is not "Unknown", then it is indeed the number of agents, i.e., $n^{\ast}=n$. Since the $\ell-b+1$ approximations of $n$ have been computed on disjoint intervals of levels, there is at least one such interval, say $\mathcal{L}_{j}$, where no leader node in the history tree has more than one child (because there can be at most $\ell$ leader branches). With the notation of Lemma 5.1, this implies that $a(\vartheta)=a(\vartheta^{\prime})$ whenever ApproxCount is called in $\mathcal{L}_{j}$. Also, since the option $x=\ell$ is tried first, the assumption $x\geq a(\vartheta)$ of Lemma 5.1 is initially satisfied. Note that ApproxCount cannot return $n^{\prime}=\texttt{"MissingNodes"}$ or $n^{\prime}=\texttt{"StrandTooShort"}$, or else $\mathcal{L}_{j}$ would not yield any approximation of $n$. Moreover, by Lemma 5.1 (ii) and by the termination condition (Line 19), if $n^{\prime}$ is not an error message while $x\geq a(\vartheta)$, then $n^{\ast}=n^{\prime}=n$. On the other hand, due to Lemma 5.1 (i), by the time $x=a(\vartheta)$ we necessarily have $n^{\prime}\neq\texttt{"WrongGuess"}$ and therefore $n^{\prime}$ is not an error message.
It remains to prove that Listing 2 actually gives an output other than "Unknown" within the claimed number of rounds; it suffices to show that it does so if it is executed at round $t^{\prime}=(\ell^{2}+\ell+1)(n-1)+1$. By Corollary 3.4, all nodes in the first $t^{\prime}-n+1=\ell(\ell+1)(n-1)+1$ levels of the history tree are contained in the vista $\mathcal{V}$ at round $t^{\prime}$. It is straightforward to prove by induction that the assumption $t^{\prime}\geq s+(\ell+2)(n-1)$ of Lemma 5.1 (iii) holds every time ApproxCount is invoked. Indeed, as long as this condition is satisfied, Lemma 5.1 (iii) implies that $n^{\prime}\neq\texttt{"MissingNodes"}$, and so "Unknown" is not returned at Line 11. Also, reasoning as in the previous paragraph, we infer that $n^{\prime}\neq\texttt{"WrongGuess"}$ by the time $x=a(\vartheta)$. Thus, within $\Delta=(\ell+1)(n-1)$ levels, either a branching leader node is found (hence $n^{\prime}=\texttt{"StrandTooShort"}$) or a new approximation of $n$ is computed (hence $n^{\prime}$ is not an error message). Every time either event occurs, $s$ is increased by at most $\Delta$ at Line 17. Thus, after $\ell-1$ (or fewer) updates of $s$, we have $s\leq(\ell-1)\Delta=t^{\prime}-(\ell+2)(n-1)-1$. Hence the condition of Lemma 5.1 (iii) holds again, and the induction goes through for at least $\ell$ steps.